Exercise 14-24 (Algo) Computing Divisional Income: Incomplete Information and Financial Ratios (LO 14-1)

As a part of an employment interview, you are given the partial income statement and selected financial ratios shown for Sneaky Pete’s, a chain of western stores. Sneaky Pete’s is organized into two divisions: Mountain and Valley. You are told that corporate overhead costs are allocated to divisions based on relative sales.

 

Required:

a. Complete the income statements for both divisions and the corporation as a whole. (Enter all values as positive value. Round your answers to 1 decimal place.)




                                       Mountain Division             Valley Division            Corporate


Sales                                                                                              

Cost of sales                                                                                  

Gross margin                                                                           

SG&A                                                                                                

Allocated corporate costs        13.8                                                                  46.0

Operating income                                                                      

Tax expense (@25%)                                                                           

After-tax income                                                                                              $127.2


Gross margin percentage       53.00%                               43.00%                46.00%Operating margin                     24.00%                               20.00%               21.20%Profit margin                            18.00%                                15.00%                15.90%
Answer #1

Answer:

                                      Mountain Division(MD)      Valley Division(VD)     Corporate


Sales                                     $240.0 (4)                         $560.0 (5)           (1) $800.0

Cost of sales                          112.8   (7)                        319.2 (15)                432.0

Gross margin                         $127.2  (6)                       $240.8 (14)               $368.0

SG&A                                       55.8  (9)                              96.6 (17)               152.4

Allocated corporate costs        13.8                                    32.2  (2)                   46.0

Operating income                   $57.6 (8)                            $112.0 (16)             $169.6

Tax expense (@25%)             14.4 (10)                             28.0  (18)                 42.4

After-tax income                     $43.2 (11)                           $84.0 (12)              $127.2


Gross margin percentage       53.00%                               43.00%             46.00% Operating margin                     24.00%                               20.00%               21.20% Profit margin                            18.00%                                15.00%                15.90%


note: (#) step number

step 1) 127.2/ 15.90%= 800          13) 560/ 800= 0.7 or 70%

2) 46.0- 13.8= 32.2                        14) 560* 43.00%= 240.8

3) 13.8/ 46.00= 0.3 or 30%            15) 560- 240.8= 319.2

4) 800* 30% (step 2)= 240.0          16) 560* 20%= 112

5) 800- 240= 560.0                         17) 240.8- 112.0- 32.2= 96.6

6) 240* 53.00%= 127.2                   18) 25%* 112=28.0

7) 240- 127.2= 112.8                      corporate column: MD+ VD

8) 240* 24.00%= 57.6

9) 127.2- 57.6- 13.8= 55.8

10) 25%* 57.6= 14.4

11) 57.6- 14.4= 43.2

12) 127.2- 43.2= 84.0